OpenStudy (anonymous):

I'm stuck on a few homework questions can someone help me figure them out? I'll fan and give medal : ]

OpenStudy (anonymous):



OpenStudy (anonymous):

Use the Triangle Sum Property to find the measure of angle y.

OpenStudy (anonymous):

@10115658 : )

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

What equation would I use for this problem?

OpenStudy (anonymous):

ok, so what do the 3 angles of a triangle add up to?

OpenStudy (anonymous):

ok, so what do the three angle of a triangle add up to?

OpenStudy (anonymous):

180

OpenStudy (anonymous):

That's right. So looking at the big triangle, we have the angles 120 and 41

OpenStudy (anonymous):

Here are my answer choices: A) 19 B) 123 C) 124 D) 111

OpenStudy (anonymous):

what's 120+41=

OpenStudy (anonymous):

161

OpenStudy (anonymous):

ok, now what's 180-161=

OpenStudy (anonymous):

oh so the answer is A) 19

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So now lets look at the smaller triangle

OpenStudy (anonymous):

so we know x is 19 and we have the other angle of 38

OpenStudy (anonymous):

can you find the other angle using the same method as before?

OpenStudy (anonymous):

I don't have to figure out x

OpenStudy (anonymous):

It was there to throw me off or over think everything and it did at first

OpenStudy (anonymous):

we haven't figured out angle y yet, that was just the first part

OpenStudy (anonymous):

we needed to find x though

OpenStudy (anonymous):

add up 19 and 38

OpenStudy (anonymous):

it said to find y

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

the problem is: Use the Triangle Sum Property to find the measure of angle y.

OpenStudy (anonymous):

yes and we needed to find the angle x in order to find the angle y. now take the angle x which is 19 and add it to 38

OpenStudy (anonymous):

180 - (19+38) =

OpenStudy (anonymous):

57

OpenStudy (anonymous):

180-57=123

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

so B is my answer not A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm still kind of confused

OpenStudy (anonymous):

So y is 123? Or 19?

OpenStudy (anonymous):

@10115658

OpenStudy (anonymous):

123

OpenStudy (anonymous):

ok